40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 18:03:47 2014, ericq, Update, Green Locking, uPDH Box Checkup uPDH_X.JPGuPDH_Y.JPGupdhTFs.pdfupdhNoises.pdfuPDH_Aug27.zip
    Reply  Thu Aug 28 17:34:20 2014, ericq, Update, Green Locking, uPDH Box Checkup TEK00006.PNG
       Reply  Fri Aug 29 17:28:07 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Fri Aug 29 23:53:35 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Tue Sep 2 20:02:06 2014, ericq, Update, Green Locking, uPDH Box Checkup 
       Reply  Wed Sep 3 00:17:22 2014, rana, Update, Green Locking, uPDH Box Checkup 
          Reply  Wed Sep 3 20:38:40 2014, ericq, Update, Green Locking, uPDH Box Checkup 
             Reply  Fri Sep 5 04:07:44 2014, ericq, Update, Green Locking, X end uPDH Box Replaced 
                Reply  Fri Sep 5 16:13:00 2014, ericq, Update, Green Locking, new X Green PDH measurements 6x
                   Reply  Sat Sep 6 01:48:54 2014, rana, Update, Green Locking, new X Green PDH measurements 
                      Reply  Sat Sep 6 14:49:12 2014, ericq, Update, Green Locking, new X Green PDH measurements xGComparison.pdfxGspectra.pdf
                         Reply  Sun Sep 7 17:06:38 2014, rana, Update, Green Locking, new X Green PDH measurements 
                         Reply  Mon Sep 8 14:14:13 2014, Jenne, Update, Green Locking, new Y Green PDH measurements 6x
Message ID: 10439     Entry time: Fri Aug 29 23:53:35 2014     In reply to: 10438     Reply to this: 10441
Author: ericq 
Type: Update 
Category: Green Locking 
Subject: uPDH Box Checkup 

I had pulled out both X and Y servo boxes for inspection, put the Y box back, soldered in a missing op amp power capacitor on the X end box, and had not yet put back the X end box yet because of the saturation issue I was looking into. Otherwise nothing was changed at the ends; I didn't open the tables at all, or touch laser/SHG settings, just unplugged the servo boxes. 

ELOG V3.1.3-